1. Compare two cups of water of the same mass one at 25ºC and the other at 85ºC. When each is contacted to a large object of temperature T, which statement is correct?

Group of answer choices

a. The heat transfer depends on the temperature of the large object.

b. The 85°C cup of water transfers more heat.

c. The 25°C cup of water transfers more heat.

d. They transfer the same total amount of heat.

2.. Imagine you transfer 5.0 J of energy to 1.0 g water and to 1.0 g ethanol, each starting at the same temperature. Which statement is true?

Cs (ethanol) = 2.4 J/g°C Cs (water) = 4.18 J/g°C

Group of answer choices

a. The ethanol will end at a higher temperature than the water.

b. Both will end at the same temperature.

c. The water will end at a higher temperature than the ethanol.

3.  You place 560 g of iron at 100°C into 56 g of water 20°C. What is the final temperature of the water (approximately - choose answer that is closest)? Cs (iron) = 0.45 J/g-°C Cs (water) = 4.18 J/g-°C

Group of answer choices

a. 60°C

b. 90°C

c. 30°C

4.. Which statement is correct about the changes in potential energy and kinetic energy when two N atoms combine to form an N2 molecule?

Group of answer choices

a. The potential energy becomes more positive and the kinetic energy increases.

b. The potential energy becomes more negative and the kinetic energy decreases.

c. The potential energy becomes more positive and the kinetic energy decreases.

d. The potential energy becomes more negative and the kinetic energy increases.
Answer #1

Question 1. Option B

Explanation: The 85°C cup of water transfers more heat because its temperature is high.

Question 2. Option A

Explanation: The ethanol will end at a higher temperature than the water because water absorbs more heat energy per oC.

Question 3. Option A

Explanation: Heat lost by iron = heat gained by the water.

i.e. 560*0.45*(100-T) = 56*4.18*(T-20)

i.e. 450 - 4.5T = 4.18T - 83.6

i.e. 8.68T = 450+83.6 = 533.6

i.e. T = 533.6/8.68 = 61.5 oC ~ 60 oC

Question 4. Option D

Explanation: When two N atoms combine to form an N2 molecule, potential energy decreases and kinetic ener